OpenStudy (anonymous):

Solve for the variable x in this equation: z = 6πxy A. z/6=x B. z/6pi=x C. z/6piy=x D. 6piyz=x

OpenStudy (anonymous):

C

OpenStudy (michele_laino):

I divide both sides of your equation by 6*pi*y, namely: \[\frac{ z }{ 6 \pi y }=\frac{ 6 \pi xy }{ 6 \pi y }=x\] so C.